3. Providing a suitable example, explain how Investment Spending (I) depends on interest  rates.

4. Based on the following equation and assumptions, answer the questions: 

For an economy, Y = C + I (here Y stands for national income, C stands for consumer spending, I stands for investment  spending) 

C = 400 + .6Y 

I = Planned Investment + Unplanned investment 

Planned investment is fixed at 500 

Planned expenditure = C + Planned Investment 

No government and closed economy

a. If Y=2500, what will be C and planned expenditure? (10) b. What is the amount of unplanned investment if Y=2500? (10) c. If Y=800, find planned expenditure. (10) 

d. Find the Y for which Y = Planned expenditure.


1
Expert's answer
2021-12-07T12:06:03-0500

(3)

Typically, higher interest rates reduce investment because higher interest rates increase the cost of borrowing and require investment to have a higher rate of return to be profitable.



As shown on the graph above, if interest rates rise from 5% to 7%, then we get a fall in the quantity of investment from 100 to 80.

If interest rates are increased, then it will tend to discourage investment because investment has a higher opportunity cost.


(4)

"Y=C+I"

"C=400+0.6Y"

Planned Investment=500

Planned Expenditure"=(400+0.6Y)+500"

(a)

If "Y=2500"

"C=400+0.6Y"

"C=400+0.6(2500)=1900"

Planned Expenditure"=C +500"

"=1900+500=2400"

(b)

"Y=2500"

"C=1900"

"Y=C+I"

"2500=1900+I"

"I=600"

Unplanned Expenditure=I- Planned Investment.

"=600-500=100"

(c)

If "Y=800,"

Planned Expenditure:

"=(400+0.6Y)+500"

"=(400+0.6\times 800)+500"

"=1380."

(d)

Y for which Y= Planned Expenditure.

"Y=(400+0.6Y)+500"

"400+500=Y-0.6Y"

"900=0.4Y"

"Y=2250"